# Constraining Super-Heavy Dark Matter with the KM3-230213A Neutrino Event

### v1

@@ IV Results
Figure 1 shows the resulting lower limits on $\tau_{\chi}$ as a function of $M_{\chi}$. The left and right panels refer to $\chi\rightarrow\nu\bar{\nu}$ and $\chi\rightarrow b\bar{b}$ channels, respectively. In general, the decay channel $\chi\rightarrow b\bar{b}$ is more strongly constrained due to its higher gamma-ray yield. Remarkably, even under our conservative assumptions, the KM3-230213A event allows us to set the most stringent constraints to date on $\tau_{\chi}$ for both decay channels considered. Specifically, in the case of $\chi\longrightarrow b\bar{b}$, we find an improvement of up to a factor of $\sim 1.5$ for $M_{\chi}\lesssim 10^{11},\rm GeV$ compared to Ref. [56], while our results are comparable at higher masses. Furthermore, for $M_{\chi}\lesssim 10^{9}\,\rm GeV$, our constraints surpass those obtained by the LHAASO collaboration [57]. Our findings reinforce previous results [11, 12, 73, 56], underscoring the power of gamma-ray observations in constraining a broad class of beyond-the-Standard-Model scenarios. Further improvements in gamma-ray upper limits at the highest energies [74] will provide complementary information and play a decisive role in probing or excluding SHDM-induced signatures even further. In the case of $\chi\rightarrow\nu\bar{\nu}$, the limits improve by up to a factor of $\sim 3$ compared to Ref. [55], and by an even larger factor relative to those reported in Ref. [5]. This enhanced sensitivity arises from a higher neutrino yield associated with this decay channel. The galactic neutrino flux, in the energy range $\gtrsim 10\,\rm PeV$, is crucial to falsify the SHDM decay hypothesis even further. Therefore, future experiments such as GRAND, RNG-0, IceCube Gen 2 and POEMMA promise unprecedented sensitivity, potentially enabling the extension of these constraints to higher energies [75, 76, 77, 78].
V Conclusions
@@
In this Letter, we have utilized the highest-energy neutrino ever detected, KM3-230213A, to establish the most stringent constraints on the SHDM decay scenario to date. Even adopting a conservative approach, we derive robust lower bounds on the dark matter lifetime, $\tau_{\chi}\gtrsim 5\cdot 10^{29}-10^{30}\rm s$ for $10^{7}\,\rm GeV\lesssim m_{\chi}\lesssim 10^{17}\,\rm GeV$, slightly varying based on the decay channel considered. Future analyses employing detailed spatial distribution templates are expected to further refine these constraints, thereby placing neutrino astronomy at the forefront of probing SHDM models.
@@
We anticipate that imminent advancements in neutrino observatories – most notably KM3NeT/ARCA and IceCube-Gen2 – with their enhanced exposure and sensitivity toward the Galactic neutrino background, complemented with upcoming ultra-high-energy gamma-ray observations, such as those from AugerPrime [74], will jointly play a crucial role in refining constraints on the SHDM decay hypothesis, potentially providing decisive evidence against hadronic decay channels or tightening constraints on leptonic channels.
@@ Acknowledgements.
Appendix A On the Direction of the KM3-230213A Event
In this section, we define the fractional visibility $F$ used in the main text to evaluate the signal number of events induced by the Galactic DM distribution. The visibility of an experiment is defined as the fraction of the day a source, in a given equatorial declination, lies within a given interval of zenith angle [79]. This limits the field of view of the experiment and it is simply dependent on the location of the detector. At ultra high energies $(E\geq 1-10\,\rm PeV)$, neutrinos can be observed specifically through down-going and horizontal events because of the Earth absorption [13, 14]. For KM3NeT/ARCA and IceCube detectors, the corresponding zenith angle range is $0^{\circ}\leq\theta\leq 100^{\circ}$ [80]. For the Pierre Auger Observatory, we consider $60^{\circ}\leq\theta\leq 95^{\circ}$ according to the range used in [15].
Fig. 2 reports the fractional visibility as a function of the declination for the experiments, evaluated through the public code available at https://zenodo.org/records/10955399 [81, 79]. The function $F(\mathrm{Declination})$ can be transformed into $F(l,b)$ through a change of coordinates, which is then inserted into Eq. 5. Remarkably, the Galactic Center is $40\%$ of the time observable in the down-going/horizontal sky for KM3NeT/ARCA. Consequently, The Galactic Center should anyway dominate in case of SHDM decay, where the flux is principally concentrated. To quantify this effect, Figs. 3 and 4 show 2D maps in galactic coordinates of the product between the visibility and the line-of-sight integral $\int ds,\rho(s,l,b)$, representing the expected contribution from Galactic dark matter decay for KM3NeT/ARCA, IceCube, and the Pierre Auger Observatory, respectively.

### v2

@@ V Results
Figure 1 shows the resulting lower limits on $\tau_{\chi}$ as a function of $M_{\chi}$. The left and right panels refer to $\chi\rightarrow\nu\bar{\nu}$ and $\chi\rightarrow b\bar{b}$ channels, respectively. In general, the decay channel $\chi\rightarrow b\bar{b}$ is more strongly constrained due to its higher gamma-ray yield. Remarkably, even under our conservative assumptions, the KM3-230213A event allows us to set the most stringent constraints to date on $\tau_{\chi}$ for both decay channels considered. Specifically, in the case of $\chi\longrightarrow b\bar{b}$, we find an improvement of up to a factor of $\sim 1.5$ for $M_{\chi}\lesssim 10^{11},\rm GeV$ compared to Ref. Das et al. (2023), while our results are comparable at higher masses. Furthermore, for $M_{\chi}\lesssim 10^{9}\,\rm GeV$, our constraints surpass those obtained by the LHAASO collaboration Cao and others (2022). Our findings reinforce previous results Abreu and others (2023b, a); Bérat et al. (2022); Das et al. (2023), underscoring the power of gamma-ray observations in constraining a broad class of beyond-the-Standard-Model scenarios. Further improvements in gamma-ray upper limits at the highest energies Castellina (2019) will provide complementary information and play a decisive role in probing or excluding SHDM-induced signatures even further. In the case of $\chi\rightarrow\nu\bar{\nu}$, the limits improve by up to a factor of $\sim 3$ compared to Ref. Chianese et al. (2021), and by an even larger factor relative to those reported in Ref. Argüelles et al. (2023). This enhanced sensitivity arises from a higher neutrino yield associated with this decay channel. The galactic neutrino flux, in the energy range $\gtrsim 10\,\rm PeV$, is crucial to falsify the SHDM decay hypothesis even further. Therefore, future experiments such as GRAND, RNG-0, IceCube Gen 2 and POEMMA promise unprecedented sensitivity, potentially enabling the extension of these constraints to higher energies Álvarez-Muñiz and others (2020); Aartsen and others (2021); Olinto and others (2021); Agarwal et al. (2025).
VI Conclusions
@@
In this manuscript, we have utilized the highest-energy neutrino ever detected, KM3-230213A, to establish the most stringent constraints on the SHDM decay scenario to date. Even adopting a conservative approach, we derive robust lower bounds on the dark matter lifetime, $\tau_{\chi}\gtrsim 5\cdot 10^{29}-10^{30}\rm s$ for $10^{7}\,\rm GeV\lesssim m_{\chi}\lesssim 10^{17}\,\rm GeV$, slightly varying based on the decay channel considered. Future analyses employing detailed spatial distribution templates are expected to further refine these constraints, thereby placing neutrino astronomy at the forefront of probing SHDM models.
@@
We anticipate that imminent advancements in neutrino observatories – most notably KM3NeT/ARCA and IceCube-Gen2 – with their enhanced exposure and sensitivity toward the Galactic neutrino background, complemented with upcoming ultra-high-energy gamma-ray observations, such as those from AugerPrime Castellina (2019), will jointly play a crucial role in refining constraints on the SHDM decay hypothesis, potentially providing decisive evidence against hadronic decay channels or tightening constraints on leptonic channels.